OpenStudy (anonymous):

help me? The cost of coffee is determined by the type of coffee beans that are mixed together. The cost of a local mix of Arabica and Robusta beans is modeled by the equation below where A = the pounds of Arabica and R = pounds of Robusta. 700A + 1200R = 1,000,000 If a mixture contains 1000 pounds of Arabica beans, how many pounds of Robusta beans should be in the mixture? R = ________________

OpenStudy (anonymous):

1714?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

yes

OpenStudy (jack1):

...@satellite73 sorry, where did we go wrong?

OpenStudy (anonymous):

i am not saying you went wrong, i am saying i don't understand it the equation given is \[700A+1200R=1,000,000\] why the cost should be fixed at 1,000,000 is beyond me, but anyway that is what it says then it says \(A=1000\) so now the equation is \[700\times 1000+1200R=1,000,000\] or \[700,000+1200R=1,000,000\] which i would solve in two steps 1) subtract \(7000,000\) to get \[1200R=300,000\] 2) then divide by \(1200\) to get \[R=250\]

OpenStudy (anonymous):

thats right..

OpenStudy (anonymous):

250 is correct

OpenStudy (jack1):

i take it all back, that's... my bad hey sorry @satellite73 and @samantha429

OpenStudy (anonymous):

lol it's okay, at least you tried :)